# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 2. Equivariant Seifert solids
Let $S$ be a periodic 2-knot in $S^{4}$. Then $S$ bounds a $\tau$-equivariant Seifert 3-manifold $Y\subset S^{4}$ such that $\operatorname{Fix}(\tau)\cap Y$ consists of an arc $A$ that is null-homologous in $Y$ rel. boundary.
The remainder of this section is devoted to the construction of $Y$.
@@
Fix an equivariant Morse function $f$ on $S^{4}$ such that its restriction to a ball around $S$ has exactly two critical points. By Proposition 1.11, we may assume that $S$ is in ordered equivariant Morse position with respect to $f$. Note that without this proposition, we would be unable to assume that the ambient Morse function induces a self-indexing Morse function on $S$, which is crucial to our construction.
It will be convenient to make use of our established conventions for coordinates for this neighborhood of $S$, as described in the introduction. In particular, we use coordinates $(x,y,z,w)$ such that $f$ is projection onto the $w$-axis and $\tau$ fixes the $(z,w)$-subspace.
First, recall that in the proof of Proposition 1.11, the critical points of the fixed set of the embedded manifold $M$ were first isotoped to be the highest maximum and lowest minimum with respect to the ambient Morse function. In the case of a 2-sphere, these are actually the only two fixed points, and so by assuming that $S$ is in ordered equivariant Morse position, we assume that the lower fixed point $p_{-}$ is the global minimum of $f|_{S}$ and that the higher, $p_{+}$, is its global maximum. In particular, the descending manifolds of each of the index 1 critical points do not contain $p_{+}$ or $p_{-}$, and so do not intersect $\operatorname{Fix}(\tau)$. Note further that apart from $p_{+}$ and $p_{-}$, all the other critical points of $K$ must come in $d$-tuples that are permuted by $\tau$. Up to translation and stretching of our coordinates, we may now assume that all the index $k$ critical values of $f|_{S}$ occur within $(k,k+\epsilon)$.
@@
We remind the reader of our notation. By $S_{w}$, we mean the preimage of $w$ under $f$ intersected with $S$, thought of as living in the whole 3-dimensional preimage of $w$. We call this preimage a $w$-slice. In this perspective, $\tau$ appears in each $w$-slice as rotation about the $z$-axis. We may construct $Y$ by specifying $Y_{w}$, defined analogously, for each $S_{w}$. In particular, for regular values $w$ of $f$, $S_{w}$ is a periodic link, and $Y_{w}$ is a Seifert surface for that link intersecting the $z$-axis in a single point. A smoothly varying family of such surfaces then stack together to form $Y$ with the desired fixed set. We will regularly switch between this perspective and explicit handle attachments to $Y$ throughout the construction. A schematic of the various handle attachments in the construction of $Y$ is shown in Figure 3 for the reader to follow along as we proceed through the construction.
[…]
Figure 3. A schematic of $Y$ showing handle attachments.
@@
$\boldsymbol{f=0.}$ Since $S$ is in ordered equivariant Morse position, we may consider all of its 0-handles first. These will be a collection of unlinked, unknotted disks. For each of these disks, we begin our construction of $Y$ by attaching a 3-dimensional index 0 right half-handle. Since all these attachments occurred in $f^{-1}(0,\epsilon)$, we see that $S_{\nicefrac{{1}}{{2}}}$ is a collection of unlinked, unknotted circles such that one of them links the $z$-axis and the rest are permuted by $\tau$, and $Y_{\nicefrac{{1}}{{2}}}$ appears as a collection of standardly embedded disks bounded by each component of $S_{\nicefrac{{1}}{{2}}}$. This is illustrated on the left of Figure 4.
@@
$\boldsymbol{f=\frac{1}{2}.}$ Now for each symmetric $d$-tuple of index 1 critical points, we consider the corresponding descending manifolds. These consist of collections of arcs passing through the index 1 critical points, possibly linking components of $S_{w}$ for $w$ below the critical value, and eventually intersecting those unknotted components. Up to vertical isotopy, these intersections can be chosen to occur in $f^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$. A schematic of this is shown on the right in Figure 4.
Figure 4. Left: the central 0-handle of $Y$, shown in multiple $w$-slices. Right: index 1 trajectories of $S$ linking slices of $S$, which would pierce the disks on the left.
@@
Because the descending manifolds may link the components of $S_{w}$ for $\frac{1}{2}+\epsilon<w<1$, we cannot simply continue the existing 0-handles of $Y$ upwards, as they would then intersect the descending manifolds. To avoid this, we first equivariantly perturb $S$ to introduce new intersections of the descending manifolds with disks bounded by $S_{\nicefrac{{1}}{{2}}}$ as necessary so that each descending arc intersects the disks in pairs of oppositely signed points. We then attach equivariant 3-dimensional interior 1-handles to $Y$ whose corresponding interior critical points have critical values in $(\frac{1}{2}-\epsilon,\frac{1}{2})$ and whose attaching spheres are vertical translates of the intersections of the index 1 boundary trajectories with the disks bounded by $S_{\nicefrac{{1}}{{2}}}$. In each $w$-slice for $w$ above $\frac{1}{2}$, we therefore see equivariant tubes attached to the disks bounded by $S_{w}$ so that the descending manifolds of the index 1 boundary critical points are engulfed by these tubes and thus do not intersect our Seifert surface in those cross-sections. This is illustrated in Figure 5.
Figure 5. A diagram of $Y_{\nicefrac{{3}}{{2}}}$, showing $S_{\nicefrac{{3}}{{2}}}$ as a periodic unlink bounding a symmetric surface. Note that the descending manifolds of the index 1 boundary critical points, depicted as bands, are prevented from intersecting the obvious disks bounded by each component by introducing tubes for them to pass through.
@@
$\boldsymbol{f=1.}$ We may now consider each index 1 critical point of $S$ as a boundary unstable index critical 1 point of $Y$, and thus for each 1-handle of $S$, we get one right half-handle attachment for $Y$. After performing the above procedure for all $d$-tuples of 1-handles in $S$, the cross-section $S_{\nicefrac{{3}}{{2}}}$ will consist again of a collection of unknotted components, one of which is a meridian for the $z$-axis. The corresponding surface $Y_{\nicefrac{{3}}{{2}}}$ will be a possibly disconnected equivariant surface of unknown genus with boundary $S_{\nicefrac{{3}}{{2}}}$.
@@
$\boldsymbol{f=\frac{3}{2}.}$ At this point, $S_{>\nicefrac{{3}}{{2}}}$ consists only of disks, but $\partial(Y_{\leq\nicefrac{{3}}{{2}}})$ is a potentially complicated equivariant surface. We would like to cap it off with equivariant 3-handles, but to do this, we must first surger it to a collection of disks. Therefore we now proceed to find a collection of equivariant compressing curves on $Y_{\nicefrac{{3}}{{2}}}$ such that the compressions do not introduce any new intersections with $\operatorname{Fix}(\tau)$.
@@
Let $\{J_{i}\}_{i\in I}$ be the set of connected components of $S_{\nicefrac{{3}}{{2}}}$, with $J_{0}$ the component that links the $z$-axis. Then the components $\{J_{i}\}_{i\in I}$ bound a pairwise disjoint collection of maximum disks $\{D_{i}\}_{i\in I}$ of $S$ such that only $D_{0}$ intersects $\operatorname{Fix}(\tau)$, and it does so transversely and exactly once. These $D_{i}$ are themselves equivariant. We may therefore equivariantly isotope copies of these disks into $f^{-1}(\frac{3}{2})$ by flowing along $\nabla f$. Call these isotoped disks $D_{i}^{\prime}$. Then equivariant $\epsilon$-neighborhoods of the disks $D_{i}^{\prime}$ for $\epsilon$ sufficiently small have boundaries that form splitting spheres $V_{i}$ separating each component of $S_{\nicefrac{{3}}{{2}}}$ from each other, and separating all $J_{i}$ except $J_{0}$ from the $z$-axis in the $\frac{3}{2}$-slice.
@@
We can now make a standard “innermost circles” argument to find equivariant compressing disks for $Y_{\nicefrac{{3}}{{2}}}$. Consider the intersections of each splitting sphere $V_{i}$ with $Y_{\nicefrac{{3}}{{2}}}$. These intersections form nested concentric circles on each sphere $V_{i}$. By compressing along these concentric disks starting with the innermost circle of intersection and working outward, we can systematically compress $Y_{\nicefrac{{3}}{{2}}}$ without introducing new intersections with the splitting spheres, as shown in Figure 6. Since these spheres are disjoint from the $z$-axis, and lie entirely in a single $w$-slice, the compressing disks are also disjoint from $\operatorname{Fix}(\tau)$, and so compressing along them does not introduce any new intersections with $\operatorname{Fix}(\tau)$. Actually performing the compressions is done by the equivariant attachment of interior 2-handles to $Y_{\leq\nicefrac{{3}}{{2}}}$, where the core disks of the handles are taken to be the compressing disks themselves. We declare that these handle attachments occur within $f^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
@@
$\boldsymbol{f=2.}$ After all such compressions are performed, $S_{\nicefrac{{3}}{{2}}+\epsilon}$ is unchanged, but $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ now consists of an equivariant collection of disjoint disks and closed components, still intersecting the $z$-axis in a single point contained inside the central disk. We further compress the closed components until they are spheres, and then fill them in with interior 3-handles. This also does not change $S_{\nicefrac{{3}}{{2}}+\epsilon}$, but surgers $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ to a collection of disks that are equivariantly isotopic to the maximum disks of $S$. Thus we may fill in the remainder of $S$ by declaring $Y_{\geq\nicefrac{{3}}{{2}}+\epsilon}$ to be trace of this isotopy. This gives $Y$, and its intersection $A$ with $\operatorname{Fix}(\tau)$ is precisely its intersection with the $w$-axis, which is a single arc as desired.
@@
It remains to show that the intersection $A$ of $Y$ with $\operatorname{Fix}(\tau)$ is null-homologous in $Y$ rel. boundary. Indeed, $A$ does not intersect the belt spheres of any 1-handles of $Y$, as all of the 1-handles were chosen to miss $\operatorname{Fix}(\tau)$. Thus $A$ is 0 in the cellular homology rel. boundary induced by the handle decomposition. ∎
We will return to our computation of the concordance group of equivariant 2-knots after a brief detour to define an invariant that will become our obstruction to equivariant sliceness.
3. An extension of the Arf invariant
3.1. Definition and related notions
The Arf invariant of a quadratic form was introduced by Arf in [Arf41] as a purely algebraic tool. Robertello gave the first definition of the Arf invariant of a knot in $S^{3}$ in [Rob65]. His paper describes three invariants of knots in $S^{3}$, then proves they are all equivalent. For our purposes, it will be helpful to recall all three. Recall that a surface $F$, possibly with boundary, properly embedded in a 4-manifold $X$ is called characteristic if $F]\in H_{2}(X,\partial X;\mathbb{Z}/2\mathbb{Z})$ is dual to the second Stiefel-Whitney class of $X$. Equivalently, for all $x\in H_{2}(X,\partial X;\mathbb{Z}/2\mathbb{Z})$, we have that $F]\cdot x\equiv x\cdot x$ mod 2, where $\cdot$ denotes intersection number rel. boundary.
Definitions 3.1 (Robertello).
Let $K$ be a knot in $S^{3}$.

### v2

@@ 3.3. The equivariant case
There exists an embedding $\tilde{\psi}^{\prime}\colon\tilde{M}\to\tilde{W}$ and an equivariant isotopy $\tilde{\Psi}\colon\tilde{M}\times I\to\tilde{W}$ between $\tilde{\psi}$ and $\tilde{\psi}^{\prime}$ such that:
  1. (1)
[…]
  2. (2)
    $\tilde{f}|_{\tilde{\psi}^{\prime}(\tilde{M})}$ has the same number and type of critical points as $\tilde{f}|_{\tilde{\psi}(\tilde{M})}$, and
  3. (3)
    $\tilde{\Psi}_{t}\equiv\tilde{\psi}$ for all $t\in I$ on a neighborhood of $\hat{A}$.
@@
Since $\tilde{f}$ and $\tilde{\psi}$ are both equivariant, we may take the quotient our whole setup by $\rho$. We then obtain an embedding $\psi\colon M\to W$ such $\psi(M)\cap\operatorname{Branch}(\rho)$ is just the image $\psi(A)$ of $\hat{A}$ after taking the quotient of $\tilde{M}$ by its periodic action. We are then in precisely the setup for Proposition 3.5, and so there exists an isotopy $\Psi$ of $M$ in $W$ rearranging $M$ to ordered Morse position with respect to the quotient Morse function $f$. Finally, we can then take the cyclic branched covers of $M$ and $W$ over their respective fixed sets and take $\tilde{\Psi}$ to be the induced isotopy upstairs, which lifts by Schwarz’s resolution of the Isotopy Lifting Conjecture [SCH80] because $\Psi$ is supported away from $A$. ∎
4. Seifert 3- and 4-manifolds
For the discussion of concordance annuli and later of periodic 2-knots in Section 6, we will need the existence of certain smooth Seifert manifolds, both in ambient dimensions 4 and 5. The Seifert 3-manifolds we construct are for spheres which represent annuli, that is, we build a Seifert 3-manifold while keeping track of its intersection with the removed sphere.
4.1. Conventions for the constructions
To aid in the following constructions, we make some parametrization conventions. Let $U$ be a 5-dimensional half-ball in $B^{5}$ intersecting the boundary in a 4-dimensional ball. We parametrize $U$ by the coordinates:
$U=\{(x,y,z,w,t)\mid x^{2}+y^{2}+z^{2}+w^{2}+t^{2}\leq C,t\geq 0\}$
for some $C\gg 1$. When we are doing constructions explicitly in $U\cap S^{4}$, we will refer to coordinates on $U$ simply as $(x,y,z,w)$. We also implicitly choose the standard metric on $U$, scaled appropriately to our choice of coordinates, which we use in our discussion of Morse theory. Submanifolds inside $U$ inherit this metric, and we will frequently differentiate between the gradient of a Morse function on the ambient space and the gradient of its restriction to the submanifold.
Finally, we use the following convention: if $f$ is a Morse function on a manifold $M$, then:
$M_{c}:=f^{-1}(c)$
for $c\in\mathbb{R}$. We also declare:
$M_{*c}:=f^{-1}(\{a\in\mathbb{R}\mid a*c\})$
where * is an element of ${<,,>,}\{$$\leq$$\geq\}$.
4.2. Seifert 3-manifolds intersecting a plane
Proposition 4.1.
Let $S$ be a smooth 2-knot in $S^{4}$, fix $R$ a smooth unknotted 2-sphere in $S^{4}$, and suppose that $S\cap R$ consists of exactly two points $p_{+}$ and $p_{-}$. Then $S$ bounds a smoothly embedded Seifert 3-manifold $Y$ in $S^{4}$ such that $Y\cap R$ is an arc $A$ bounded by $p_{+}$ and $p_{-}$ which is null-homologous in $Y$ rel. boundary.
Proof.
@@ 4.2. Seifert 3-manifolds intersecting a plane
Fix a Morse function $f$ on $S^{4}$ such that its restriction to a ball around $S$ has exactly two critical points. By Proposition 3.5, we may assume that $S$ is in ordered Morse position with respect to $f$. As noted in the previous subsection, we use coordinates $(x,y,z,w)$ such that $f$ is projection onto the $w$-axis and $R$ appears as the $(z,w)$-subspace.
@@
Recall that by $S_{w}$, we mean the preimage of $w$ under $f$ intersected with $S$, thought of as living in the whole 3-dimensional preimage of $w$. We call this preimage a $w$-slice. We may construct $Y$ by specifying $Y_{w}$, defined analogously, for each $S_{w}$. In particular, for regular values $w$ of $f$, $S_{w}$ is a link which does not intersect the $z$-axis, and $Y_{w}$ is a Seifert surface for that link intersecting the $z$-axis in a single point. A smoothly varying family of such surfaces then stack together to form $Y$ with the desired fixed set. We will regularly switch between this perspective and explicit handle attachments to $Y$ throughout the construction.
@@
$\boldsymbol{f=0.}$ Since $S$ is in ordered Morse position, we may consider all of its 0-handles first. These will be a collection of unlinked, unknotted disks. For each of these disks, we begin our construction of $Y$ by attaching a 3-dimensional index 0 right half-handle. Since all these attachments occurred in $f^{-1}(0,\epsilon)$, we see that $S_{\nicefrac{{1}}{{2}}}$ is a collection of unlinked, unknotted circles such that one of them links the $z$-axis, and $Y_{\nicefrac{{1}}{{2}}}$ appears as a collection of standardly embedded disks bounded by each component of $S_{\nicefrac{{1}}{{2}}}$.
@@
$\boldsymbol{f=\frac{1}{2}.}$ Now for each index 1 critical point, we consider the corresponding descending manifold. These consist of collections of arcs passing through the index 1 critical points, possibly linking components of $S_{w}$ for $w$ below the critical value, and eventually intersecting those unknotted components. Up to vertical isotopy, these intersections can be chosen to occur in $f^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$.
@@
Because the descending manifolds of the index 1 critical points of $S$ may link the components of $S_{w}$ for $\frac{1}{2}+\epsilon<w<1$, we cannot simply continue filling in the 0-handles $S$ with disks past $w=\frac{1}{2}$, as they would then intersect the descending manifolds. To avoid this, we first perturb $S$ to introduce new intersections of the descending manifolds with disks bounded by $S_{\nicefrac{{1}}{{2}}}$ as necessary so that each descending arc intersects the disks in pairs of oppositely signed points. We then attach 3-dimensional interior 1-handles to $Y$ whose corresponding interior critical points have critical values in $(\frac{1}{2}-\epsilon,\frac{1}{2})$ and whose attaching spheres are vertical translates of the intersections of the index 1 boundary trajectories with the disks bounded by $S_{\nicefrac{{1}}{{2}}}$. In each $w$-slice for $w$ above $\frac{1}{2}$, we therefore see tubes attached to the disks bounded by $S_{w}$ so that the descending manifolds of the index 1 boundary critical points are engulfed by these tubes and thus do not intersect our Seifert surface in those cross-sections.
Figure 10. Tubes attached to $Y_{\nicefrac{{1}}{{2}}}$ removing intersections with descending manifolds as a result of interior 1-handle attachments.
@@
$\boldsymbol{f=1.}$ We may now consider each index 1 critical point of $S$ as a boundary unstable index critical 1 point of $Y$, and thus for each 1-handle of $S$, we attach one right half-handle to $Y$. After performing the above procedure for all 1-handles in $S$, the cross-section $S_{\nicefrac{{3}}{{2}}}$ will consist again of a collection of unknotted components, one of which is a meridian for the $z$-axis. The corresponding surface $Y_{\nicefrac{{3}}{{2}}}$ will be a possibly disconnected surface of unknown genus with boundary $S_{\nicefrac{{3}}{{2}}}$.
@@
$\boldsymbol{f=\frac{3}{2}.}$ At this point, $S_{>\nicefrac{{3}}{{2}}}$ consists only of disks, but $Y_{\nicefrac{{3}}{{2}}}$ is a potentially complicated surface. We would like to cap it off with 3-handles, but to do this, we must first surger it to a collection of disks. Therefore we now proceed to find a collection of compressing curves on $Y_{\nicefrac{{3}}{{2}}}$ such that the compressions do not introduce any new intersections with $R$.
@@
Let $\{J_{i}\}_{i\in I}$ be the set of connected components of $S_{\nicefrac{{3}}{{2}}}$, with $J_{0}$ the component that links the $z$-axis. Then the components $\{J_{i}\}_{i\in I}$ bound a pairwise disjoint collection of maximum disks $\{D_{i}\}_{i\in I}$ of $S$ such that only $D_{0}$ intersects $R$, and it does so transversely and exactly once. We may therefore isotope copies of these disks into $f^{-1}(\frac{3}{2})$ by flowing along $\nabla f$. Call these isotoped disks $D_{i}^{\prime}$. Then $\epsilon$-neighborhoods of the disks $D_{i}^{\prime}$ for $\epsilon$ sufficiently small have boundaries that form splitting spheres $V_{i}$ separating each component of $S_{\nicefrac{{3}}{{2}}}$ from each other, and separating all $J_{i}$ except $J_{0}$ from the $z$-axis in the $\frac{3}{2}$-slice.
@@
We can now make a standard “innermost circles” argument to find compressing disks for $Y_{\nicefrac{{3}}{{2}}}$. Consider the intersections of each splitting sphere $V_{i}$ with $Y_{\nicefrac{{3}}{{2}}}$. These intersections form nested concentric circles on each sphere $V_{i}$. By compressing along these concentric disks starting with the innermost circle of intersection and working outward, we can systematically compress $Y_{\nicefrac{{3}}{{2}}}$ without introducing new intersections with the splitting spheres, as shown in Figure 11. Since these spheres are disjoint from the $z$-axis, and lie entirely in a single $w$-slice, the compressing disks are also disjoint from $R$, and so compressing along them does not introduce any new intersections with $R$. Actually performing the compressions is done by attaching interior 2-handles to $Y_{\leq\nicefrac{{3}}{{2}}}$, where the core disks of the handles are taken to be the compressing disks themselves. We declare that these handle attachments occur within $f^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
@@
$\boldsymbol{f=2.}$ After all such compressions are performed, $S_{\nicefrac{{3}}{{2}}+\epsilon}$ is unchanged, but $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ now consists of a collection of disjoint disks and closed components, still intersecting the $z$-axis in a single point contained inside a single disk. We further compress the closed components until they are spheres, and then fill them in with interior 3-handles. This also does not change $S_{\nicefrac{{3}}{{2}}+\epsilon}$, but surgers $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ to a collection of disks that are isotopic to the maximum disks of $S$. Thus we may fill in the remainder of $S$ by declaring $Y_{\geq\nicefrac{{3}}{{2}}+\epsilon}$ to be trace of this isotopy. This gives $Y$, and its intersection with $R$ is precisely its intersection with the $w$-axis, which is a single arc as desired.
@@
It remains to show that the intersection $A$ of $Y$ with $R$ is null-homologous in $Y$ rel. boundary. Indeed, $A$ does not intersect the belt spheres of any 1-handles of $Y$, as all of the 1-handles were chosen to miss $R$. Thus $A$ represents the 0 class in the cellular homology rel. boundary induced by the handle decomposition. ∎
4.3. Seifert 4-manifolds intersecting a 3-ball
We now prove a similar statement to Proposition 4.1, but one dimension up. Note that for this construction, there are extra Morse-theoretic conditions that were not present in the previous. It is unknown at this time whether these can be removed, and this question is closely tied to the discussion in Remark 1.14.
Proposition 4.2.
Fix $Q$ a smoothly unknotted 3-sphere in $S^{5}$. Let $Y$ be a closed 3-manifold smoothly embedded in $S^{5}$, and suppose that $Y\cap Q$ is a null-homologous knot $K$ in $Y$. Suppose that there exists a Morse function $f$ on $S^{5}$ such that $Y$ is in ordered Morse position with respect to $f$, $f|_{K}$ has exactly two critical points, and away from these two critical points, $K$ is fully vertical with respect to $f|_{Y}$. Then $Y$ bounds a smooth Seifert 4-manifold $X$ smoothly embedded in $S^{5}$ such that $X\cap Q$ is a smoothly embedded disk $D$ bounded by $K$ which is null-homologous in $X$ rel. boundary.
Proof.
We follow roughly the same strategy as in the proof of Proposition 4.1. That is, we build $X$ handle by handle according to a handle decomposition of $Y$ coming from $f$. We do this, as before, by finding fillings for slices of $Y$, which themselves are surfaces. To better visualize these surfaces, we parametrize a neighborhood of $Y$ so that $Q$ appears locally as the $(z,w,t)$-subspace, and so that $f$ is just projection to the $t$ coordinate. Note that the full verticality condition on $K$ ensures that the lowest minimum (maximum) of $Y$ is necessarily also the minimum (resp. maximum) of $K$, and so every $t$-slice has at least one intersection point with $Q$ in it. The assumption that $K$ is fully vertical with respect to $f|_{Y}$ away from two critical points implies that no $t$-slice of $Y$ intersects $Q$ in more than two points. Thus in each $t$-slice apart from those containing the very lowest minimum and the very highest maximum, $Y$ appears as a surface in 4-space intersecting the $(z,w)$-subspace in two points. Our goal is therefore to construct smoothly varying 3-dimensional fillings $X_{t}$ of these surfaces $Y_{t}$ such that each of these slices intersects the $(z,w)$-subspace in a single smoothly varying arc. These arcs stacked on top of each other will then form a disk $D$, which will be $X\cap Q$ as desired.
$\boldsymbol{f=0.}$ Since $Y$ is in ordered Morse position, we can assume that all of its 0-handles occur in $0\leq t<\epsilon$, with the lowest minimum intersecting $Q$ in an unknotted arc. We may begin constructing $X$ by filling in these 0-handles with 4-dimensional index 0 right half-handles, such that the aforementioned arc bounds a disk given by a straight line in the $(z,w)$-plane connecting each intersection point at each level. This completes $X$ and $D$ up to $t=\frac{1}{2}$.
$\boldsymbol{f=\frac{1}{2}.}$ At this point, $X_{\nicefrac{{1}}{{2}}}$ is a collection of 3-balls in ambient 4-space. As in the previous case, we want to fill the 1-handles of $Y$ with 4-dimensional right half-handles. However, the trajectories from the index 1 critical points of $Y$ may link slices of the 0-handles of $Y$. As before, we consider these trajectories intersecting the 0-handles of $Y$ in $f^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$. To avoid these intersections, we preemptively attach interior 1-handles to $X$ at $t=\frac{1}{2}-\epsilon$ along the trajectories to create 3-dimensional tubes for these trajectories to pass through. Since the trajectories from the 1-handles of $Y$ do not intersect $Q$, and all interior 1-handles are attached parallel to these trajectories, we can assume that none of the 1-handles of $X$ intersect $Q$. Thus we may continue constructing $D$ by taking the straight line in the $(z,w)$-plane connecting $Y_{t}\cap Q$ in each $t$-slice.
$\boldsymbol{f=1.}$ After these interior 1-handles are attached, we may proceed to attach 4-dimensional index 1 right half-handles to fill each 1-handle of $Y$. Again, since none of the 1-handles of $Y$ intersect $Q$, these half-handles do not either, and we can continue constructing $D$ by taking the straight line in the $(z,w)$-plane connecting the two points of $Y_{t}\cap Q$ in each $t$-slice.
$\boldsymbol{f=\frac{3}{2}.}$ Following the same general strategy, we now examine the descending manifolds of the 3-dimensional index 2 critical points at $t=2$. Because $Y\cap Q$ is precisely $K$, we know none of these descending manifolds intersect $Q$, as all such intersections are already accounted for by our assumption that $K$ has precisely two critical points with respect to the ambient Morse function. Thus in theory we may safely attach right half-handles to $X$ for each 2-handle of $Y$ without fear of introducing new components of $X\cap Q$. However, the index 2 descending manifolds may intersect interior 1-handles of $X$ that we have already attached. As before, suppose that all of these intersections occur within the preimage $f^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
To avoid these intersections, consider an index 2 critical point of $Y$. Note that the generic intersection between a 2-handle core and a 4-dimensional 1-handle in $B^{5}$ is a disk in the 2-handle core. Thus we can use this disk as a guide to preemptively attach 4-dimensional interior 2-handles to compress the 4-dimensional 1-handles to avoid intersecting the 3-dimensional 2-handle. This is completely analogous to the preemptive 1-handles we attached to make way for the 1-handles of $Y$. Note further that because the disks we use as guides for attaching our 4-dimensional 2-handles are already embedded in $B^{5}$ as the cores of existing 3-dimensional 2-handles, they have trivial normal bundles. Thus from the perspective of $X$, the compressions are achieved by attaching interior 2-handles, with framings given by a choice of trivial disk sub-bundle of the normal bundle. Furthermore, since none of the 2-handles of $Y$ intersect $Q$, we may arrange so that the parallel 2-handles of $X$ do not intersect $Q$ either.
$\boldsymbol{f=2.}$ Once we have performed these compressions at $t=\frac{3}{2}-\epsilon$, we may safely proceed upwards in $t$ and fill the 3-dimensional 2-handles of $Y$ with 4-dimensional index 2 half-handles at $t=2$.
$\boldsymbol{f=\frac{5}{2}.}$ At this point, for $t\in(2+\epsilon,\frac{5}{2}-\epsilon)$, $Y_{t}$ appears as an unlink of unknotted spheres intersecting the $(z,w)$-plane in exactly two points on a single component, and $X_{t}$ appears as a connected 3-manifold with boundary $Y_{t}$ intersecting the $(z,w)$-plane in a single straight line. Now every component of $Y_{\nicefrac{{5}}{{2}}-\epsilon}$ bounds a 3-ball embedded in $B^{5}_{\geq 2+\epsilon}$ given by the descending manifolds of each maximum of $f|_{Y}$. As in the 3-manifold case, we may isotope these 3-ball caps straight down into the $t=\frac{5}{2}-\epsilon$-slice, using their boundary unknotted 2-spheres as guides near the boundary.
As in the lower-dimensional case, the boundaries of tubular neighborhoods of each of these 3-balls then form splitting spheres for $Y_{\nicefrac{{5}}{{2}}-\epsilon}$. These splitting spheres intersect $X_{\nicefrac{{5}}{{2}}-\epsilon}$ in a collection of 3-balls, which we may then use as compressing 3-balls for $X_{\nicefrac{{5}}{{2}}-\epsilon}$. We can then compress $X$ along these 3-balls by attaching interior 3-handles to $X_{\nicefrac{{5}}{{2}}}$ and filling in closed components.
$\boldsymbol{f=3.}$ In fact, because the splitting spheres we found were the boundaries of small neighborhoods of balls bounded by components of $Y_{\nicefrac{{5}}{{2}}}$, each component of $X_{\nicefrac{{5}}{{2}}}$ with boundary is itself a 3-ball. We may subsequently fill the remaining closed components with equivariant 4-manifold fillings. Finally, we trace the isotopy of the 3-ball caps up to the maxima of $Y$ to find 4-balls that fill these maxima, with the central 4-ball intersecting $Q$ in a disk. This completes the construction of $X$.
We now consider the disk $D$ that we have constructed as $\operatorname{Fix}(\rho)\cap X$. As in the case of $A$ in the proof of Proposition 6.2, $D$ does not intersect the belt spheres of any handles used in the construction of $X$, and thus represents the 0 class in the relative homology of $X$. ∎
4.4. Computing the Arf invariant using Seifert 4-manifolds
The purpose of Proposition 4.2 is to show that the Arf invariant defined in Section 2 can be computed using a Seifert 4-manifold:
Proposition 4.3.
Let $Y,Q$, and $K$ be as in Proposition 4.2. Then $\operatorname{Arf}(K,Y,\mathfrak{s}_{0})=0$, where $\mathfrak{s}_{0}$ is the restriction of the unique spin structure on $B^{5}$ to $Y$.
Proof.
The key observation is that the construction in Proposition 6.3 gives a characteristic filling of $(K,Y,\mathfrak{s}_{0})$. Since $X$ embeds in $S^{5}$, $X$ is necessarily spin, and so since $[D]=0$ in the relative homology of $X$, $D$ is characteristic in $X$. The fact that $X$ is spin also means that $\sigma(X)=8\mu(Y,\mathfrak{s}_{0})$. In fact, spin structure on $Y$ induced by inclusion into $B^{5}$ extends over $X$ because $X$ is also embedded in $B^{5}$, and so the canonical spin characterization of $D$ in $X$ extends the canonical spin characterization of $K$ in $Y$ as they are both induced by the same ambient spin structure. Finally, since $K$ is null-homologous in $Y$, it has a unique Seifert framing in $Y$, and so the fact that $D$ is null-homologous rel. boundary implies that its normal Euler number relative to the Seifert framing is 0. We may therefore compute:
$\operatorname{Arf}(K,Y,\mathfrak{s}_{0})=\frac{D\cdot D-\sigma(X)}{8}+\mu(Y,\mathfrak{s})=\frac{0}{8}-\frac{\sigma(X)}{8}+\frac{\sigma(X)}{8}=0.$
5. Constructing and obstructing concordances
Following Proposition 4.3, we can now define our concordance invariant for Montesinos annuli.
5.1. The concordance invariant for annuli
Definition 5.1.
Let $N$ be a Montesinos annulus in $S^{1}\times S^{3}$, and let $S$ be its corresponding sphere in $S^{4}$ with removed sphere $R$. Let $Y$ be an Seifert 3-manifold for $S$ with $Y\cap R=A$ a null-homologous arc in $Y$. We define the invariant $\alpha$ of $N$ to be:
$\alpha(N):=\operatorname{Arf}(A,Y,\mathfrak{s}_{0})\in\mathbb{Z}/2\mathbb{Z},$
where $\mathfrak{s}_{0}$ denotes the spin structure on $Y$ induced by the unique spin structure on $B^{5}$ via inclusion.
The goal of this section is to prove the following classification theorem.
Theorem 5.2.
There exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive relative concordance invariant $\alpha$ for Montesinos annuli $N$ in $S^{1}\times B^{3}$ such that: